OpenStudy (anonymous):

Compute a second-order Taylor series expansion around point (a,b) = (0,0) of a function f(x,y)=e^x*log(1+y)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Taylor_series The example at the end of the page is exactly your problem

OpenStudy (amistre64):

a taylor series of multivariables correlates to the partial derivatives

OpenStudy (amistre64):

Fx is the term for x Fxy is the term for xy and so on

OpenStudy (anonymous):

Another way to this is to write \[e^x =1+x+\frac{x^2}{2}+\frac{x^3}{6 }+O\left(x^4\right)\\ \ln( 1+ y)=y-\frac{y^2}{2}+\frac{y^3}{3}+ O\left(y^4\right)\\ e^x \ln(1+y)=\left ( 1+x+\frac{x^2}{2}+\frac{x^3}{6 }+O\left(x^4\right) \right) \left ( y-\frac{y^2}{2}+\frac{y^3}{3}+ O\left(y^4\right) \right) \] Multiply out and you can find some few terms of your series expansion.